RC meeting and Run Plan, Wednesday Oct. 17

From HallCWiki
Jump to: navigation, search

Status

  • Beam Quality Bottom line:
    • Yesterday, accelerator had issues getting a good beam back after RF recovery. They finally settled on 50uA as maximum stable current, and we requested 45uA to limit the SHMS singles rate to below 700kHz in the current setting.
    • Today around noon accelerator went down to fix some more Hall B related issues, as well as do some work on the accelerator.


Issues: Resolved

Issues: Requiring attention

Ongoing

Requests

  • Hall D needs opportunistic 1shift of HallC down for beam studies, not pressing right now (needs to be done before end-of-run)
    • Might be possible when need access for deuterium target.
  • Dave Meekins needs controlled access during work on deuterium target Monday starting 8am
    • It will take ~4 hours
  • Swapping out aerogel will be Tuesday during RF recovery

Plans - Physics

  • As of right now,(4-30pm) we have collected 38% of the statistics goal for LH2 on Q2=5.5, W=3.02, theta_Shms=6.49.
  • We were originally planning to move to the next kinematic setting around the start of swing, but we will most likely push this to 9PM.
    • After that we will try to get a minimum amount of data in the last two remaining settings. You can find the tentative run plan for the Kaon LT End Game here:

Run Plan for Kaon LT 10.6GeV End Game